OpenStudy (anonymous):

AP CALCULUS BC Solve for x: 3/(x+3) _< 2/(x-2) (the sign means smaller than or equal to) I found (-infinity, -3) U (2, 12] Is that correct?

OpenStudy (anonymous):

so \(x\ne2 \text{, and }x\ne-3\). then\[ 3\left( x-2 \right) \le 2\left( x+3 \right) \Rightarrow 3\left( x-2 \right)-2\left( x+3 \right) \le 0 \]\[\Rightarrow x-12\le 0 \Rightarrow x \le 12 \Rightarrow \left(-\infty , -3\right) \cup \left(-3, 2\right)\cup \left( 2, 12 \right) \]

OpenStudy (anonymous):

you're right, the middle doesn't work.

OpenStudy (anonymous):

yeah the middle interval goes -/- which is positive

OpenStudy (anonymous):

glad i got the right answer, thanks!

OpenStudy (anonymous):

great job!!!